# Processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Inert Higgs Doublet Models and Two Higgs Doublet Models

### v2

@@ 4.1.2 Enhancement factors
The enhancement factors given in Eq. 42 are examined in the IHDM. In Fig. 9, the factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are scanned in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2}$. In the following scatter plots, singly charged Higgs masses are varied from $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $-200$ GeV $\leq\mu_{2}\leq 200$ GeV. Furthermore, we fix $\lambda_{2}=0.8$ and $M_{H}=150$ GeV for all cases. The data is generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (all left panel scatter-plots) GeV and at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (all right panel scatter-plots).
The factors for $hh$ productions in the IHDM are first analyzed. Since the cross-sections for $hh$ productions are enhanced around the peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}$. Therefore, it is not surprised in finding that $\mu_{hh}^{\textrm{IHDM}}$ becomes largest at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). Around these peaks, the data indicates that the enhancement factors tend to about $\sim 1.5$ in the limit of $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$. Since the contributions of charged Higgs in the loop being small when $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$ (due to the fact that the couplings of $hH^{\pm}H^{\mp}$, $hhH^{\pm}H^{\mp}$ tend to zero in this limit). It is found that the enhancement factors can reach to factor of $6$ (for $500$ GeV of CoM) and factor of $13$ (for $1000$ GeV of CoM) around the peaks. Beyond the peaks, we observe that $1\leq\mu_{hh}^{\textrm{IHDM}}\leq 2$.
For the enhancement factors of $HH$ productions in the IHDM, we also find $\mu_{HH}^{\textrm{IHDM}}$ becomes largest contributions at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). It is attentive to realize that $\mu_{HH}^{\textrm{IHDM}}$ have different behavior in comparison with $\mu_{hh}^{\textrm{IHDM}}$. At the $500$ GeV of CoM, the factors develop to the peak and then are decreased rapidly beyond the peak. However, they grow up with the charged Higgs masses in the above regions of $M_{H^{\pm}}\geq\sim 300$ GeV. Because there aren’t couplings $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ only depends on $\lambda_{2}$. Therefore, in the high regions of charged Higgs masses, the factors $\mu_{HH}^{\textrm{IHDM}}$ are increased with $M_{H^{\pm}}$.
@@
In Fig. 10, the enhancement factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are generated in the space of $M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2}$. The charged Higgs masses are varied as $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the scatter plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for all left panel plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all right panel plots). For the factors $\mu_{hh}^{\textrm{IHDM}}$ (as shown in all the above scatter plots), both the couplings $hH^{\pm}H^{\mp}$ and $hhH^{\pm}H^{\mp}$ are independent of $\lambda_{2}$. As a result, the factors only depend on $M_{H^{\pm}}$. For the factors $\mu_{HH}^{\textrm{IHDM}}$ (as shown in all the below scatter plots), it is found that the quadratic-coupling $HHH^{\pm}H^{\mp}$ depend on $\lambda_{2}$. As a result, the factors depend strongly on $\lambda_{2}$ and $M_{H^{\pm}}$. These massive contributions are mainly from the charged Higgs exchanging in the box diagrams.
@@
Figure 10: The enhancement factors are scanned over the parameter space of $M_{H^{\pm}},\lambda_{2}$. Charged Higgs masses are in $70$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above-Figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below-Figures), correspondingly.
@@ 4.2 THDM
The phenomenological results for the production processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the THDM are analysed in the following subsection.
@@ 4.2.1 Production cross-sections
We first inspect the data in the case of $M^{2}=0$. One notices that $\hat{\sigma}_{HH}$ become largest in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq\sim 450$ GeV and they are decreased rapidly in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV. Moreover, $\hat{\sigma}_{hh}$ in the SM and the THDM are dominant in the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV contrasted to the ones for $\gamma\gamma\rightarrow hH,\leavevmode\nobreak\ HH$ in the THDM. Among the mentioned cross-sections, the $hh$ production in the THDM is largest in this case.
@@
Fermionphobic limit
The fermionphobic limit is studied in which the mixing angle is taken as $\alpha=\pm\pi/2$. For a typical example, we select $\alpha=+\pi/2$ in the following plots. Since, we have already checked that top quark propagating in the loop is dominant contributions versus other fermions. It is enough to take into account top quark in the loop for the present calculations. It means that the cross-sections are only contributed from $W$ boson and scalar particles in the loop when we consider the fermionphobic limit. Subsequently, we can examine the comparative sizes among these contributions. In Fig. 12 (for $hh$ production), Fig. 13 (for $hH$ production), Fig. 14 (for $HH$ production), the corresponding cross-sections for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM together with $\hat{\sigma}_{hh}$ in the SM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$, are analysed in the fermionphobic limit. The CoM energies are varied $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We select $M_{H^{\pm}}=300$ GeV and apply $t_{\beta}=3$ (blue line), $t_{\beta}=5$ (green line), $t_{\beta}=7$ (red line), respectively. Moreover, we fix $M_{H}=150$ GeV for all cases. In further, the results are presented at $M^{2}=0$ GeV2 (for all the left Figures) and at $M^{2}=+200^{2}$ GeV2 (for all the right Figures).
The $hh$ productions in the SM and the THDM at $M^{2}=0$ GeV2 are first analysed. We notice that the cross-sections depend slightly on $t_{\beta}$ in the regions below the peak $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=600$ GeV. In the regions above the peak, it is found that the cross-sections are more sensitive to $t_{\beta}$. For the case of $M^{2}=200^{2}$ GeV2, the cross-sections are proportional to $t_{\beta}$. Around the peak regions, the cross-sections are enhanced by charged Higgs loop.
$\hat{\sigma}_{hh}(M^{2}=0)$ [pb] $\hat{\sigma}_{hh}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 12: Cross-sections for $\gamma\gamma\rightarrow hh$ in the SM and the THDM as functions $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the fermionphobic limit. We show the data for $M^{2}=0$ GeV2 on the left panel and for $M^{2}=+200^{2}$ GeV2 on the right panel.
As mentioned in above, the cross-sections for $hH$ productions in the THDM are suppressed due to the softly breaking of the $Z_{2}$-symmetry. It is explainable for cross-sections for $hH$ productions are much smaller than the corresponding ones for $hh$ productions in the SM. However, at the peak of $\sqrt{\hat{s}_{\gamma\gamma}}=2M_{H^{\pm}}=600$ GeV, the cross-sections are enhanced and can reach to order of $\hat{\sigma}_{hh}$ in the SM. At $M^{2}=0$, $\hat{\sigma}_{\gamma\gamma\rightarrow hH}$ are more sensitive to $t_{\beta}$ in all range of CoM. Another case of $M^{2}=+200^{2}$ GeV2, the cross-sections are also more sensitive to $t_{\beta}$ in the regions below the peak $\sim 600$ GeV. But they are nearly proportional to $t_{\beta}$ beyond the peak.
$\hat{\sigma}_{hH}(M^{2}=0)$ [pb] $\hat{\sigma}_{hH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 13: Cross-sections for $\gamma\gamma\rightarrow hH$ in the SM and the THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ in the fermionphobic limit. In the plots, we vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We present the data for $M^{2}=0$ GeV2 on the left panel and for $M^{2}=+200^{2}$ GeV2 on the right panel.
For all cases of $M^{2}$ in $HH$ productions, it is interested in finding that the cross-sections are larger than $\hat{\sigma}_{hh}$ in the SM and they are proportional to $1/t_{\beta}$ for all range of CoM. We also observe that cross-sections are enhanced around the peak $\sim 2M_{H^{\pm}}=600$ GeV. The dominant contributions are from the singly charged Higgs exchanging in the loop.
$\hat{\sigma}_{HH}(M^{2}=0)$ [pb] $\hat{\sigma}_{HH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 14: Cross-sections for $\gamma\gamma\rightarrow HH$ in the SM and THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ in the fermionphobic limit. In the plots, we vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We show the data for $M^{2}=0$ GeV2 on the left panel and for $M^{2}=+200^{2}$ GeV2 on the right panel.
Decoupling limit
[…]
$\hat{\sigma}_{hh}(M^{2}=0)$ [pb] $\hat{\sigma}_{hh}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 15: Cross-sections for $\gamma\gamma\rightarrow hh$ in the SM and THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the decoupling limit.
[…]
$\hat{\sigma}_{hH}(M^{2}=0)$ [pb] $\hat{\sigma}_{hH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 16: Cross-sections for $\gamma\gamma\rightarrow hH$ in the THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the decoupling limit.
$\hat{\sigma}_{HH}(M^{2}=0)$ [pb] $\hat{\sigma}_{HH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 17: Cross-sections for $\gamma\gamma\rightarrow HH$ in the THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the decoupling limit.
@@ 4.2.2 Enhancement factors
We pay attention to investigate the enhancement factors defined in Eq. 42 for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM. The factors scanned over parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$ are first studied in this subsection. Two scenarios for $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are examined in detail. In Figs. 18,  19, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2. Moreover, we vary $100$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and set $2\leq t_{\beta}\leq 10$ in the following plots. The factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ are generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for all above scatter-plots) and examined at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all below scatter-plots). In the left panel, we show for the enhancement factors for $hh$ productions. In the middle (right) panel, the enhancement factors for $hH,\leavevmode\nobreak\ (HH)$ productions are presented, respectively.
In Fig. 18, the first scenario for $c_{\beta-\alpha}>0$ is explored. In this scenario, we take $c_{\beta-\alpha}=+0.2$ for an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, correspondingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, $\mu_{hh}^{\textrm{THDM}}$ change from $1$ to $1.5$ for all range of $M_{H^{\pm}}$. The values of $\mu_{hh}^{\textrm{THDM}}$ are enhanced around the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=250$ GeV. Predominantly, the factors are proportional to $t_{\beta}^{-1}$ in this case. Interestingly, we observe that the factors $\mu_{hH}^{\textrm{THDM}}$ change from 00 to $0.5$ for all range of $M_{H^{\pm}}$. The suppressed values of $\mu_{hH}^{\textrm{THDM}}$ are expected as explained in previous paragraphs due to the $Z_{2}$-symmetry. The $\mu_{hH}^{\textrm{THDM}}$ are the same behavior as $\mu_{hh}^{\textrm{THDM}}$ which they are inversely proportional to $t_{\beta}$. In the other hand, the enhancement factors for $HH$ productions in the THDM are strongly dependent of the charged Higgs mass but change slightly with $t_{\beta}$. In all range of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ are from $0.3$ to $1.2$.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, the factors $\mu_{hh}^{\textrm{THDM}}$ become biggest at the peak at $M_{H}^{\pm}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ change from $1.2$ to $1.8$. Beyond the peak, the factors are changed from $1.0$ to $1.2$ in all range of $M_{H^{\pm}}$. It is stress that $\mu_{hh}^{\textrm{THDM}}$ slightly change with $t_{\beta}$. For $HH$ productions, $\mu_{HH}^{\textrm{THDM}}$ are varied from $0.4$ to $2.5$ around the peak (at $M_{H^{\pm}}=500$ GeV) regions. It is realized that $\mu_{HH}^{\textrm{THDM}}$ slightly change with $t_{\beta}$. Otherwise, $\mu_{hH}^{\textrm{THDM}}$ are much smaller than $1$ and are inversely proportional to $t_{\beta}$.
@@
Figure 18: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$. In the plots, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2 and we take $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for the above-plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below plots), respectively.
@@
Another scenario for $c_{\beta-\alpha}<0$ is considered for examining how are the factors effect by setting different sign of $c_{\beta-\alpha}$ in this work. In Fig. 19, we take $c_{\beta-\alpha}=-0.2$ for an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, it is excited in observing that $\mu_{hh}^{\textrm{THDM}}$ are different behavior in comparison with previous scenario. At this CoM energy, the factors $\mu_{hh}^{\textrm{THDM}}$ can reach to $1.5$ in the low region of $M_{H^{\pm}}<200$ GeV. They then are decreased to around $0.9$ when $M_{H^{\pm}}>200$ GeV. There isn’t peak of the factors observed in this scenario. Because the contributions of singly charged Higgs exchanging in the one-loop triangle diagrams may cancel out with the ones one-loop box diagrams in this scenario. Surprisingly, we find that the factors $\mu_{hh}^{\textrm{THDM}}$ are proportional to $t_{\beta}$ in this scenario. For the $hH$ productions, the factors are suppressed and they are in the range of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ]. They are sensitive with $t_{\beta}^{-1}$ in all range of charged Higgs mass. In $HH$ productions, it is found that the factors develop to the peak around $M_{H^{\pm}}=500$ GeV. They reach to factor $2.5$ around the peak and they are in the ranges of [0.5,1.7][$\sim$ 0.5$\sim$ 1.7][ ∼ 0.5 , ∼ 1.7 ] beyond the peak regions. In all range of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ are proportional to $t_{\beta}^{-1}$ in this scenario.

### v3

@@ 3.1.2 Enhancement factors
The enhancement factors given in Eq. 28 are examined in the IHDM. In Fig. 2, the factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are scanned in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2}$. In the following scatter plots, singly charged Higgs masses are varied from $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $-200$ GeV $\leq\mu_{2}\leq 200$ GeV. Furthermore, we fix $\lambda_{2}=0.8$ and $M_{H}=150$ GeV for all cases. The data are generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (all left panel scatter-plots) GeV and at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (all right panel scatter-plots).
The factors for $hh$ production in the IHDM are first analyzed. Since the cross-sections for $hh$ production are enhanced around the peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}$, it is not surprising to find that $\mu_{hh}^{\textrm{IHDM}}$ becomes largest at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). Around these peaks, the data indicate that the enhancement factors tend to about $\sim 1.5$ in the limit of $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$, since the contributions of the charged Higgs in the loop are small when $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$ (due to the fact that the couplings of $hH^{\pm}H^{\mp}$ and $hhH^{\pm}H^{\mp}$ tend to zero in this limit). It is found that the enhancement factors can reach a factor of $6$ (for $500$ GeV of CoM) and a factor of $13$ (for $1000$ GeV of CoM) around the peaks. Beyond the peaks, we observe that $1\leq\mu_{hh}^{\textrm{IHDM}}\leq 2$.
For the enhancement factors of $HH$ production in the IHDM, we also find that $\mu_{HH}^{\textrm{IHDM}}$ becomes largest at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). It is important to realize that $\mu_{HH}^{\textrm{IHDM}}$ has different behavior in comparison with $\mu_{hh}^{\textrm{IHDM}}$. At $500$ GeV of CoM, the factors rise to the peak and then decrease rapidly beyond the peak. However, they grow with the charged Higgs masses in the above regions of $M_{H^{\pm}}\geq\sim 300$ GeV. Because there are no couplings of $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ only depends on $\lambda_{2}$, the factors $\mu_{HH}^{\textrm{IHDM}}$ increase with $M_{H^{\pm}}$ in the high regions of charged Higgs masses.
@@
In Fig. 3, the enhancement factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are generated in the space of $M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2}$. The charged Higgs masses are varied as $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the scatter plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for all left panel plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all right panel plots). For the factors $\mu_{hh}^{\textrm{IHDM}}$ (as shown in all the above scatter plots), both the couplings $hH^{\pm}H^{\mp}$ and $hhH^{\pm}H^{\mp}$ are independent of $\lambda_{2}$. As a result, the factors only depend on $M_{H^{\pm}}$. For the factors $\mu_{HH}^{\textrm{IHDM}}$ (as shown in all the below scatter plots), it is found that the quadratic-coupling $HHH^{\pm}H^{\mp}$ depends on $\lambda_{2}$. As a result, the factors depend strongly on $\lambda_{2}$ and $M_{H^{\pm}}$. These massive contributions are mainly from the charged Higgs exchanging in the box diagrams.
@@
Figure 3: The enhancement factors are scanned over the parameter space of $M_{H^{\pm}},\lambda_{2}$. Charged Higgs masses are in $70$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below figures), correspondingly.
@@ 3.2 THDM
The phenomenological results for the production processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the THDM are analyzed in the following subsection.
@@ 3.2.1 Production cross-sections
We inspect the data in the case of $M^{2}=0$. One notices that $\hat{\sigma}_{HH}$ becomes largest in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq\sim 450$ GeV and they decrease rapidly in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV. Moreover, $\hat{\sigma}_{hh}$ in the SM and the THDM are dominant in the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV contrasted to the ones for $\gamma\gamma\rightarrow hH,\leavevmode\nobreak\ HH$ in the THDM. Among the mentioned cross-sections, the $hh$ production in the THDM is largest in this case.
@@ 3.2.2 Enhancement factors
We focus on investigating the enhancement factors defined in Eq. 28 for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM. The factors are first studied over the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$ in this subsection. Two scenarios for $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are examined in detail. In Figs. 5,  6, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2. Moreover, we vary $100$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and set $2\leq t_{\beta}\leq 10$ in the following plots. The factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ are generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for all the above scatter-plots) and examined at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all the below scatter-plots). In the left panel, we show the enhancement factors for $hh$ production. In the middle (right) panel, the enhancement factors for $hH$ and $(HH)$ production are presented, respectively.
In Fig. 5, the first scenario for $c_{\beta-\alpha}>0$ is explored. In this scenario, we take $c_{\beta-\alpha}=+0.2$ as an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, correspondingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, $\mu_{hh}^{\textrm{THDM}}$ changes from $1$ to $1.5$ for the entire range of $M_{H^{\pm}}$. The values of $\mu_{hh}^{\textrm{THDM}}$ are enhanced around the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=250$ GeV. Predominantly, the factors are proportional to $t_{\beta}^{-1}$ in this case. Interestingly, we observe that the factors $\mu_{hH}^{\textrm{THDM}}$ change from 00 to $0.5$ for the entire range of $M_{H^{\pm}}$. The suppressed values of $\mu_{hH}^{\textrm{THDM}}$ are expected, as explained in the previous paragraph, due to the $Z_{2}$-symmetry. The $\mu_{hH}^{\textrm{THDM}}$ exhibit the same behavior as $\mu_{hh}^{\textrm{THDM}}$, in that they are inversely proportional to $t_{\beta}$. On the other hand, the enhancement factors for $HH$ productions in the THDM are strongly dependent on the charged Higgs mass but change slightly with $t_{\beta}$. In the entire range of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ range from $0.3$ to $1.2$.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, the factors $\mu_{hh}^{\textrm{THDM}}$ become the largest at the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ changes from $1.2$ to $1.8$. Beyond the peak, the factors change from $1.0$ to $1.2$ in the entire range of $M_{H^{\pm}}$. It is important to note that $\mu_{hh}^{\textrm{THDM}}$ slightly changes with $t_{\beta}$. For $HH$ productions, $\mu_{HH}^{\textrm{THDM}}$ varies from $0.4$ to $2.5$ around the peak (at $M_{H^{\pm}}=500$ GeV) regions. It is noted that $\mu_{HH}^{\textrm{THDM}}$ slightly changes with $t_{\beta}$. Otherwise, $\mu_{hH}^{\textrm{THDM}}$ is much smaller than $1$ and is inversely proportional to $t_{\beta}$.
@@
Figure 5: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$. In the plots, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2 and take $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above plots) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below plots), respectively.
@@
Another scenario for $c_{\beta-\alpha}<0$ is considered for examining how the factors are affected by setting a different sign of $c_{\beta-\alpha}$ in this work. In Fig. 6, we take $c_{\beta-\alpha}=-0.2$ for an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, it is interesting to observe that $\mu_{hh}^{\textrm{THDM}}$ shows different behavior in comparison with the previous scenario. At this CoM energy, the factors $\mu_{hh}^{\textrm{THDM}}$ can reach up to $1.5$ in the low region of $M_{H^{\pm}}<200$ GeV. They then decrease to around $0.9$ when $M_{H^{\pm}}>200$ GeV. There is no peak of the factors observed in this scenario because the contributions of singly charged Higgs exchanging in the one-loop triangle diagrams may cancel out with the ones from the one-loop box diagrams in this scenario. Surprisingly, we find that the factors $\mu_{hh}^{\textrm{THDM}}$ are proportional to $t_{\beta}$ in this scenario. For the $hH$ productions, the factors are suppressed and they are in the range of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ]. They are sensitive to $t_{\beta}^{-1}$ in all ranges of charged Higgs mass. In $HH$ productions, it is found that the factors develop to the peak around $M_{H^{\pm}}=500$ GeV. They reach a factor of $2.5$ around the peak and they are in the ranges of [0.5,1.7][$\sim$ 0.5$\sim$ 1.7][ ∼ 0.5 , ∼ 1.7 ] beyond the peak regions. In all ranges of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ are proportional to $t_{\beta}^{-1}$ in this scenario.